OpenStudy (anonymous):

how do you solve this problem... 5x + 2 (4+x) = 10

jigglypuff314 (jigglypuff314):

distributive property is like a(b+c) = ab + ac so 2(4+x) = 8 + 2x so 5x + 2(4+x) = 10 5x + 8 + 2x = 10 make sense so far? :)

jigglypuff314 (jigglypuff314):

then you can add 5x + 2x = 7x so 5x + 8 + 2x = 10 becomes 7x + 8 = 10 then subtract 8 from both sides -8 -8 7x = 2 then divide both sides by 2

OpenStudy (anonymous):

@jigglypuff314 divide both sides by 7 7x/7=2/7 x=2/7

jigglypuff314 (jigglypuff314):

oh right, sorry ;)

OpenStudy (anonymous):

:)